# KeyB2: Selecting Key Blocks is Also Important for Long Document Ranking with Large Language Models

### v1

@@ Abstract.
The rapid development of large language models (LLMs) like Llama has significantly advanced information retrieval (IR) systems. However, using LLMs for long documents, as in RankLLaMA, remains challenging due to computational complexity, especially concerning input token length. Furthermore, the internal mechanisms of LLMs during ranking are still not fully understood. In this paper, we first explore the internal workings of LLMs during relevance judgement and identify that specific attention heads play a crucial role in aligning relevant tokens. This observation inspires us to revisit the block pre-ranking strategy used in KeyB, which remains state-of-the-art (SOTA) on the TREC 2019 DL document ranking dataset. Building on these insights, we develop KeyB2, an advanced long document IR approach that integrates block pre-ranking with the performance of LLMs. KeyB2 efficiently identifies and processes the most relevant blocks, reducing computational costs and improving ranking effectiveness. Additionally, we introduce a new bi-encoder block matching strategy for KeyB2. Comprehensive experiments on long-document datasets, including TREC 2019 DL, Robust04, and MLDR-zh, show that KeyB2 outperforms baselines like RankLLaMA and KeyB by reducing reranking time and GPU memory usage while enhancing retrieval performance, achieving new SOTA results on TREC 2019 DL with higher NDCG@10 and MAP scores.
@@ 1. Introduction
The rapid development of large language models (LLMs) such as GPT and Llama has transformed diverse natural language processing (NLP) tasks, including information retrieval (IR), text generation, and comprehension. These models have demonstrated remarkable capabilities in capturing deep semantic relationships within text, leading to significant advancements in IR systems. Cross-encoder models, such as RankLLaMA (Ma et al., 2024), have been particularly effective at evaluating the relevance of queries and passages by encoding entire input sequences jointly. Despite these advancements, processing long documents with LLMs remains challenging, primarily due to the quadratic complexity of self-attention mechanisms, leading to high computational and memory demands. Furthermore, there remains room for improvement in the effectiveness of directly using LLMs.
@@
What occurs within LLMs, such as RankLLaMA, during relevance judgements in IR tasks, and how do attention mechanisms influence the judgement of query and document content?
@@
How can the KeyB approach be upgraded using LLMs, and what are the best strategies for designing effective block selection methods?
@@
Can the integration of block selection strategies with LLMs improve both the efficiency and effectiveness of long document ranking compared to models that process entire documents? Is block selection still crucial in the era of advanced LLMs for long document retrieval, and does it provide an advantage over directly processing entire documents?
We validate our approach with extensive experiments across several long-document datasets, including TREC 2019 DL document ranking task, Robust04, and MLDR-zh. Our findings show that KeyB2 significantly outperforms baseline models, including traditional and LLM-based approaches, in terms of retrieval accuracy. Besides, KeyB2 models have higher computational efficiency compared with LLM based IR approach RankLLaMA. The results highlight that combining LLMs with intelligent block selection can optimize long document ranking while maintaining or improving retrieval performance.
The structure of this paper is as follows: Section 2 reviews related work in IR, LLMs, and block selection methods. Section 3 provides a detailed analysis of attention mechanisms within LLMs and their relevance to document retrieval. Section 4 introduces the KeyB2 approach, while Section 5 outlines the experimental setup, and presents the results of effectiveness and efficiency with analysis. Lastly Section 6 concludes the paper.
@@ 2.1. From Neural-IR Models to Pre-trained Language Models for Information Retrieval
Transformer-based models, such as BERT (Devlin et al., 2019), have been particularly effective. They leverage attention mechanisms, pre-trained on large scale textual data, enabling a deeper understanding of the semantics embedded within queries and documents. The success of BERT in various NLP tasks, including IR, can be attributed to its capacity to obtain contextualized understanding that account for the surrounding text. It can be classified into three categories of BERT-based IR methods: interaction-based approach, representation-based approach and late-interaction based approach. The first kind is also called cross-encoder, where the query and passage are concatenated and inputted to the BERT model. Nogueira and Cho (2019) proposed using BERT as a cross-encoder for passage re-ranking by fine-tuning the model, achieving significantly higher results than earlier neural-IR models.
@@
ColBERT (Khattab and Zaharia, 2020), a representative approach of the third category, introduces a late interaction architecture that independently encodes each token in the query and the document using BERT, and then employ a cheap yet powerful interaction step that models the fine-grained similarity. Later, ColBERTv2 (Santhanam et al., 2022) is proposed that employs a simple combination of distillation from a cross-encoder and hard-negative mining to boost quality. Then a residual compression mechanism is used to reduce the storage space of the representations.
@@ 2.3. Reranking with Large Language Models
In conclusion, most of these approaches focus on passage retrieval tasks rather than on long-document ranking. RankLLaMA experiment on TREC 2019 DL document ranking task, but the results are lower than KeyB with BERT model. In this paper, we naturally follow the standard RankLLaMA approach and want to see if selecting key block approach is also important in the LLM era and can enhance the results.
3. Attention Mechanisms in LLMs for Information Retrieval
@@
To shed light on this, we propose analyzing the attention heatmaps of the rankLLaMA model, which has been fine-tuned on the MSMARCO passage ranking dataset111available at: https://huggingface.co/castorini/rankLlama-v1-7b-lora-passage. A similar investigation by (Clark et al., 2019), even though on a different architecture with less layers and attention heads reveals that a significant portion of BERT’s attention is focused on the delimiter token [SEP], which often attends broadly over entire sentences. In their work, they also observe that certain attention heads align well with linguistic features such as syntax and coreference. However, recent LLMs are uni-directional and decoder-only, and may display a different behavior regarding how tokens attend to each other. We used the BertViz tool (Vig, 2019) to explore these attention patterns in the decoder-only LLM RankLLaMA model when processing various query-passage pairs.
@@
(c) More attention to current tokens
@@
(e) ¡s¿ token exhibits broad attention
(f) ¡/s¿ token exhibits broad attention
(g) Broad attention
(h) ¡s¿ token exhibits broad attention
(i) ¡/s¿ token exhibits broad attention
@@
(d) The overall attention map
(e) The attention map for a specific token
(f) The attention map for a specific token
Figure 2. Example attention map of two heads, that potentially are capturing the relevance information.
@@
(d) The overall attention map
(e) The attention map for a specific token
(f) The attention map for a specific token
Figure 3. Attention map of two heads that deal with relevant and irrelevant information. To help visualization, top 3 images are sharpened.
@@
(d) The attention map for a specific token
@@
Figure 4. An example attention map of a head, that captures more relevant information.
3.1. Analysis
@@
We begin with a simple and arbitrary text pair: the query text is ”where is Paris”, and the document text is ”Paris is a city in France”. So the input text to the LLM is ”¡s¿query: where is Paris¡s¿document: Paris is a city in France¡/s¿”. Our findings are summarized in the following points:
@@
  • Attention to delimiter, current and broad tokens: Similar to (Clark et al., 2019) in BERT, we observe that a substantial portion of attention is directed towards the delimiter token ¡s¿, e.g., Fig. 1(a). Besides, we also observe a small amount of attention is directed towards the current token (Fig. 1(b), Fig. 1(c)). The authors in (Clark et al., 2019) speculate that attention over the ¡s¿ delimiter tokens might be used as a sort of “no-op”, and we conjecture that the attention over the beginning ¡s¿ and current tokens is also ”no-op” for IR scenario. We also observe that some attention heads attend broadly over all tokens (Fig. 1(d), Fig. 1(g)), which is consistent with (Clark et al., 2019).
    The delimiter tokens, as ¡s¿ and ¡/s¿, are important in PLMs and LLMs as, respectively the CLS token, separation token or the last delimiter token, are used to separate the input or compute the retrieval status value (RSV) of a document to a given query. It is thus important to examine what happens with these delimiter tokens, and in particular to the last delimiter for Llama-based models. We can see in Fig. 1(e) that the second token ¡s¿ attends broadly over the query tokens, and on Fig. 1(f) that the token ¡/s¿ attends broadly over both the query and the document tokens. This suggests that these delimiters capture the overall representations of, respectively, the query and of the query-document pair. In Fig. 1(h) and Fig. 1(i), they also perform similarly but attend more to salient tokens, in the query for the second ¡s¿ token and in the document for the token ¡/s¿. As the last delimiter ¡/s¿, at the basis of the RSV calculation in RankLLaMA, also exhibits broad attention, it can eventually be viewed as a coarse-grained representation of the query-document pair dedicated to IR.
  • Attention between relevant tokens: Interestingly, when examining other attention heads, we find that several heads seem to capture relations between relevant tokens in the query and document. As shown in Fig. 2, for Attention Head 25 in Layer 1, the token “Paris” in the document attends primarily to “Paris” in the query, while the token “France” in the document attends to “Paris” in both the query and document. For Attention Head 25 in Layer 8, tokens in the document tend to attend to relevant tokens in the query. These illustrate that specific attention heads are adept at identifying relevant matches between query terms and document content, which is essential for accurate relevance judgement.
  • Attention with irrelevant information: To test what happens between irrelevant tokens, we introduce irrelevant information, such as “An apple is a round fruit”, into the document. As shown in Fig. 3, the attention head 25 in layer 8 suppresses the irrelevant content, maintaining strong attention on the relevant tokens “Paris” and “France.” This head is trying to discard the noise (e.g., apple) and focuses on query-relevant portions. However, head 25 in layer 1, which displays broad attentions over all tokens, shows that the attention connections can be more contrasted. Fig. 3(b) shows that the token “France” attends more to Paris than to other tokens, and that the last token ¡/s¿ attends to both relevant tokens, “Paris” and “city”, as well as irrelevant tokens like “apple”, “is”, “a”, although weakly. These irrelevant tokens will inevitably result in noisy information in the representation of the last token “¡/s¿”, which is the basic building block for computing the relevance score for RankLLaMA.
  • Dynamic adjustment with additional relevant information: We further analyze how the model handles additional relevant information by expanding the document to include another sentence about Paris, as shown in Fig. 4 (to be compared with Fig. 3). The attention map shows that the attention head tries to adjust its focus to cover both instances of relevant content. This is similar to the way humans construct relevance judgments for long documents (Wu et al., 2007), by assessing the relevance of blocks within documents.
3.2. Discussion
These insights reveals part of the internal processes behind RankLLaMA. The attention mechanisms in Llama models, as explored in this section, show similarities to the way humans assess the relevance of documents, by selectively focusing on relevant information while filtering out irrelevant content. For example, given a query like ”Where is Paris?” in a passage such as ”Paris is a city in France. An apple is a round fruit,” humans naturally prioritize the tokens ”Paris” and ”France,” while ignoring irrelevant parts like ”apple.” Similarly, Llama models are trying to concentrate their attention on query-relevant terms, as seen in Fig. 2, Fig. 3 and Fig. 4. However, certain attention heads attend broadly, as shown in Fig. 3(a) and Fig. 3(c), and the last token ¡/s¿ weakly attends to irrelevant tokens, resulting in noisy representations and potentially erroneous relevance scores.
Although recent LLMs can handle input sizes from thousands to hundreds of thousands of tokens, their computational complexity still depends on the input token count. Similar to the Transformer or BERT architectures, recent LLMs like Llama2 or Llama3 also use self-attention mechanism. The difference is recent LLMs use the decoder-only architecture with unidirectional attention. Llama models use Grouped Query Attention (GQA) (Ainslie et al., 2023): instead of each attention head having its own key (K) and value (V) matrices, multiple heads are grouped together, and each group shares the same key and value vectors. This mechanism is designed to reduce memory and computational overhead, particularly in large models like Llama and Mistral (Jiang et al., 2023), by allowing several query heads to reuse the same key-value pair within a group. This helps speed up inference while maintaining the benefits of multi-head attention. Recent LLMs also use KV Cache (Pope et al., 2023) to speed up inference, but the KV Cache still faces memory bound issue. Besides, the cache does not apply to RankLLaMA. Indeed, the KV Cache technique caches the Key (K) and Value (V) matrices from previous tokens, enabling the model to reuse these stored results when generating new tokens. This avoids redundant computations and significantly improves efficiency. However, the RankLLaMA model only use the ¡/s¿ token to infer the relevance once and does not need to generate more tokens. Based on the above analyses, the complexity of the RankLLaMA model remains quadratic relative to the input token count.
The block selection strategy for IR (Li et al., 2023a) mimics human filtering by choosing the most relevant passages or blocks from long documents, reducing computational load and improving retrieval performance by focusing on critical content and removing noisy contents. Pre-ranking blocks enhances both the efficiency and accuracy of relevance judgements with BERT. Similarly, for LLMs used in IR, such as RankLLaMA, a pre-ranking strategy for passages or blocks may help models avoid irrelevant tokens and the noisy representations they may lead to, and improve both efficiency and effectiveness in long document retrieval tasks. This is the question we are addressing now.
4. Method: KeyB2

### v2

@@ Abstract.
The emergence of large language models (LLMs) such as Llama has significantly advanced neural information retrieval (IR). However, applying LLMs to long document reranking remains computationally expensive and may be ineffective. Moreover, the internal behavior of LLMs during document relevance judgment is still underexplored. In this paper, we begin with an in-depth analysis of decoder-only LLM attention patterns and find that several attention heads consistently align with relevance signals, yet this alignment deteriorates as irrelevant content increases. Motivated by this observation, we revisit and extend the block selection paradigm, introducing KeyB2, a scalable reranking framework that combines block pre-selection with powerful decoder-only LLMs. KeyB2 generalizes the selection stage to support BM25, cross-encoder, and bi-encoder, and adapts LLM to compute fine-grained relevance scores. We further introduce a new bi-encoder strategy that performs strongly and efficiently. Extensive experiments on TREC DL 2019/2023 document task, Robust04, and MLDR-zh demonstrate that KeyB2 outperforms baselines including RankLLaMA, RankLLaMA-MaxP/AvgP, and KeyB, achieving new state-of-the-art (SOTA) results on TREC DL 2019 document reranking task. In addition, KeyB2 reduces reranking latency compared with RankLLaMA by over 83% and memory usage by over 74%, positioning it as a practical and effective solution for long document ranking with LLMs.
@@ 1. Introduction
The rapid development of LLMs such as GPT and Llama has transformed diverse natural language processing (NLP) tasks, including IR, text generation, and comprehension. These models have demonstrated remarkable capabilities in capturing deep semantic relationships within text, leading to significant advancements in IR systems. Cross-encoder models, such as RankLLaMA (Ma et al., 2024), have been particularly effective at evaluating the relevance of queries and passages by encoding entire input sequences jointly. Despite these advancements, processing long documents with LLMs remains challenging, primarily due to the quadratic complexity of self-attention mechanisms, leading to high computational and memory demands. Furthermore, there remains room for improvement in the effectiveness of directly using LLMs.
@@
Interpretability. How do large-language-model rerankers allocate attention between query and document tokens during relevance estimation, and do the resulting patterns suggest that relevance cues are concentrated in a limited set of blocks?
@@
Ranking Effectiveness. Does pairing an LLM with an explicit block selector (KeyB2) improve long-document ranking effectiveness over the the KeyB approach and full-document LLM reranker (RankLLaMA)?
@@
Selector Design. Which block-selection strategy—lexical (BM25), neural bi-encoder, or neural cross-encoder—achieves the best balance between selection quality and computational overhead when integrated with an LLM backbone?
RQ4:
Computational Efficiency. To what extent does block selection reduce inference latency and GPU memory compared with full-document LLM rerankers, and where do KeyB2 variants lie on the accuracy–efficiency Pareto frontier?
Our contributions are summarized as follows:
  • We provide an in-depth attention analysis of LLM rerankers (RankLLaMA), opening the black box and showing that full-document LLMs struggle with attention dispersion in long-document retrieval tasks, especially in noisy contexts, motivating the need for explicit block selection.
  • We propose KeyB2, a generalization of the KeyB framework that integrates block selection into LLM-based reranking. KeyB2 extends the original KeyB design by adopting stronger LLaMA2 and LLaMA3 backbones and supporting flexible block selectors (BM25, bi-encoder, cross-encoder).
  • We conduct comprehensive experiments on four datasets (TREC DL19/23 document reranking, Robust04, MLDR-zh), covering in-domain, cross-domain, and bi-lingual scenarios. KeyB2 consistently outperforms prior baselines, including full-document RankLLaMA and previous KeyB models, achieving new SOTA on TREC 2019 DL document reranking task.
  • We systematically analyze the computational efficiency of KeyB2, demonstrating that block selection allows substantial reductions in inference latency and GPU memory usage while maintaining or improving ranking accuracy.
The structure of this paper is as follows: Section 2 reviews related work in LLM-based information retrieval, including recent developments in reranking and block selection. Section 3 conducts a comprehensive attention-based analysis of RankLLaMA to investigate whether it internally captures relevance between queries and long documents, thereby addressing RQ1. Section 4 introduces the proposed KeyB2 approach, including its architecture and differences from prior KeyB methods. Section 5 and Section 6 present the experimental setup and results, analyzing retrieval effectiveness (RQ2), robustness under noisy conditions (RQ3), and computational efficiency (RQ4). Finally, Section 7 concludes the paper and outlines future directions.
@@ 2.1. From Neural-IR Models to Pre-trained Language Models for Information Retrieval
Transformer-based models, such as BERT (Devlin et al., 2019), have been particularly effective. They leverage attention mechanisms, pre-trained on large scale textual data, enabling a deeper understanding of the semantics embedded within queries and documents. The success of BERT in various NLP tasks, including IR, can be attributed to its capacity to obtain contextualized understanding that account for the surrounding text. It can be classified into three categories of BERT-based IR methods: interaction-based approach, representation-based approach and late-interaction based approach. The first kind is also called cross-encoder, where the query and passage are concatenated and inputted to the PLM or LLM. Nogueira and Cho (2019) proposed using BERT as a cross-encoder for passage re-ranking by fine-tuning the model, achieving significantly higher results than earlier neural-IR models.
@@
ColBERT (Khattab and Zaharia, 2020), a representative approach of the third category, introduces a late interaction architecture that independently encodes each token in the query and the document using BERT, and then employs a cheap yet powerful interaction step to model fine-grained similarity. Later, ColBERTv2 (Santhanam et al., 2022b) is proposed, which combines distillation from a cross-encoder with hard-negative mining to boost effectiveness. It also introduces a residual compression mechanism to reduce the storage cost of token-level representations. To further improve efficiency, PLAID (Santhanam et al., 2022a) builds on ColBERTv2 with a performance-optimized engine that leverages centroid interaction and pruning to reduce search latency by up to 45× on CPU and 7× on GPU, while maintaining high retrieval quality.
@@ 2.3. Reranking with Large Language Models
In conclusion, most of these approaches focus on passage retrieval tasks rather than on long-document ranking. RankLLaMA experiment on TREC 2019 DL document reranking task, but the results are lower than KeyB with BERT model. In this paper, we naturally follow the standard RankLLaMA approach and want to see if selecting key block approach is also important in the LLM era and can enhance the results.
3. Attention Analysis: Why Block Selection Still Matters
@@
(c) Broad attention
@@
(d) The overall attention map (noise)
(e) The attention map for a specific token (noise)
(f) The attention map for a specific token (noise)
Figure 2. Example attention map of layer 1 head 23, that potentially is capturing the relevance information.
@@
(d) The overall attention map (noise)
(e) The attention map for a specific token (noise)
(f) The attention map for a specific token (noise)
Figure 3. Example attention map of layer 8 head 25, that potentially is capturing the relevance information.
@@
(d) The overall attention map (noise)
(e) The attention map for a specific token (noise)
(f) The attention map for a specific token (noise)
@@
Figure 4. Example attention map of layer 31 head 24, that potentially is capturing the relevance information.
3.1. Attention Heatmaps of Specific Examples
To shed light on this and answer RQ1, we propose first analyzing the attention heatmaps of the RankLLaMA model with several specific examples, which has been fine-tuned on the MSMARCO passage ranking dataset111available at: https://huggingface.co/castorini/rankLlama-v1-7b-lora-passage. Clark et al. (2019) propose a similar investigation for BERT and find a significant portion of BERT’s attention is focused on the delimiter token, and certain attention heads align well with linguistic features such as syntax and coreference. However, recent LLMs are uni-directional and decoder-only, especially for the IR focused LLMs, and may display a different behavior regarding how tokens attend to each other. We used the BertViz tool (Vig, 2019) to explore these attention patterns in the decoder-only LLM RankLLaMA when processing various query-passage pairs.
@@ 3.1. Attention Heatmaps of Specific Examples
We begin with a simple and arbitrary text pair: the query text is ”where is Paris”, and the document text is ”Paris is a city in France”. With the format ”query: [query] document: [document]”, after Llama2 tokenizer, the input tokens are: ‘¡s¿’, ‘_query’, ‘:’, ‘_where’, ‘_is’, ‘_Paris’, ‘_document’, ‘:’, ‘_Paris’, ‘_is’, ‘_a’, ‘_City’, ‘_in’, ‘_France’. We check several attention heads which are shown in Fig. 1, Fig. 2, Fig. 3 and Fig. 4. To help showing weak attention weights, the last three figures are sharpened. Our findings are summarized in the following points:
@@
  • Attention to delimiter, current and broad tokens: Similar to (Clark et al., 2019) in BERT, we observe that a substantial portion of attention is directed towards the delimiter token ¡s¿, e.g., Fig. 1(a). Clark et al. (2019) speculate that attention over the ¡s¿ delimiter tokens might be used as a sort of “no-op”, and we conjecture that the attention over the beginning ¡s¿ is also ”no-op” for IR scenario. Besides, we also observe a small amount of attention is directed towards the current token (Fig. 1(b)), and that some attention heads attend broadly over all tokens (Fig. 1(c), Fig. 1(d)), which is consistent with (Clark et al., 2019).
  • Attention focusing on relevant tokens: Beyond delimiter-focused or self-focused heads, we observe several attention heads that directly capture semantic relations between relevant tokens in the query and document. In particular, as shown in each above three subfigures in Fig. 2, Fig. 3, and Fig. 4, these relevance-focused heads consistently highlight key cross-token alignments: for head 23 in layer 1, tokens such as “Paris” and “France” in the document primarily attend to corresponding relevant tokens in both query and document segments; for head 25 in layer 8, tokens “Paris” and “France” focus strongly on the query token “Paris”; for head 24 in layer 31, similar alignment between query and document relevance tokens is also preserved. These patterns suggest that a subset of attention heads are able to capture fine-grained token-level relevance signals, forming direct associations between query intents and document content—an essential mechanism underlying accurate relevance estimation.
  • Attention with irrelevant information: To investigate the effect of irrelevant tokens, we insert unrelated content such as “An apple is a fruit” into the document. This setting is illustrated in the three subfigures in Fig. 2, Fig. 3, and Fig. 4. The attention head 25 in layer 8 (Fig. 3(e)) effectively suppresses the irrelevant content, maintaining strong attention on the relevant query token “Paris.” This head appears to filter out the noise (e.g., apple) and prioritize query-relevant content.
    In contrast, head 23 in layer 1 (Fig. 2(e)) continues to exhibit broad attention across all tokens. Although the attention may be weak, it reveals that the head does not clearly distinguish between relevant and irrelevant content. Similarly, head 24 in layer 31 (Fig. 4(e)) shows that the token “Paris” in the document attends not only to the query token “Paris” but also weakly to irrelevant tokens such as “apple,” “is,” and “a.”
    Importantly, except for head 25 in layer 8, the introduction of irrelevant tokens leads to a noticeable reduction in the attention weight toward the query token “Paris.” For instance, in Fig. 2(f) and Fig. 4(f), the attention from the final document token “France” to the query token “Paris” becomes weaker after inserting irrelevant information.
    To support this observation quantitatively, Table 1 presents the measured attention scores from “France” to “Paris” in the query across the three heads. All heads exhibit decreased attention scores following the insertion of irrelevant tokens. Notably, while the decline for Layer 8 Head 25 is mild, the other two heads show substantial drops. This degradation may impair the model’s ability to correctly assess the relevance between document and query, which is particularly critical in information retrieval scenarios. These irrelevant tokens will inevitably result in noisy information in the representation of the last token, which is the basic building block for computing the relevance score for RankLLaMA.
Table 1. The attention weights between “France” to “Paris” in the query when inserting noise tokens.
Layer 1 Head 23 (Fig 2) Layer 8 Head 25 (Fig 3) Layer 31 Head 24 (Fig 4)
no noise tokens 0.2097 0.6500 0.1296
with noise tokens 0.1154 (↓44.97%) 0.6492 (↓0.12%) 0.0858 (↓33.80%)
(a) The overall attention heatmaps for relevant documents (without inserting noise)
(b) The overall attention heatmaps when inserting noise before relevant documents
(c) The overall attention heatmaps when inserting noise after relevant documents
Figure 5. The overall attention heatmaps across samples, comparing with inserting noise tokens.
3.2. Aggregated Attention Heatmaps Across Examples
The previous section illustrates attention behaviours on individual examples. To investigate whether these patterns generalize across real-world long documents, we conduct a dataset-level aggregation analysis. Specifically, we sample 500 query–relevant document pairs from the TREC DL19 document ranking’s development set (MS MARCO) (Craswell et al., 2020). The RankLLaMA model fine-tuned for document-level reranking [^2] https://huggingface.co/castorini/rankllama-v1-7b-lora-doc is used for evaluation. All documents are truncated to 1200 tokens.
For each example, we compute the average attention scores from document tokens to query tokens across all heads in every layer. The attention scores are then averaged across the sampled set to produce an overall heatmap summarizing attention behaviour across layers and heads.
To assess robustness against irrelevant context, we also create two additional test settings. For each query–relevant document pair, we randomly sample a negative document, extract its first 800 tokens as noise, and insert the noise either before or after the relevant document. The aggregated attention heatmaps under these two noise injection settings are similarly computed. The resulting heatmaps for the clean and noisy cases are shown in Fig. 5.
3.2.1. Findings
As shown in Fig. 5(a), certain heads—primarily located in the beginning and middle layers, as well as several heads in the final layers—exhibit strong attention from document tokens to query tokens, aligning with relevance-focused behaviour observed earlier. However, when noise is inserted, both Fig. 5(b) and Fig. 5(c) display a general attenuation of attention scores, indicating that irrelevant content weakens the focus on relevant tokens. Notably, inserting noise after the relevant content appears to cause greater attention dispersion, likely because the model processes text autoregressively from left to right.
These aggregated patterns confirm that a fraction of attention heads identify relevance signals, and that irrelevant text can dilute these signals, motivating selective block extraction before LLM reranking.
3.3. Quantitative Analysis with Attention-Relevance Alignment Score (ARAS) and Positive Correlation Rate (PCR)
To complement the above qualitative attention observations and provide a more systematic and quantitative understanding of how LLM rerankers behave (addressing RQ1), we propose two evaluation metrics: the Attention-Relevance Alignment Score (ARAS) and the Positive Correlation Rate (PCR). These metrics aim to measure how well the model’s attention aligns with true relevance signals in long documents.
3.3.1. Metric Definitions
  • Attention weight per chunk: Given a query-document pair $(q,d)$, we first segment the document $d$ into $M$ non-overlapping chunks $\{C_{1},C_{2},\dots,C_{M}\}$, where each chunk contains a fixed number of tokens. For a given attention head and layer, let $\mathbf{A}\in\mathbb{R}^{L\times L}$ denote the attention matrix for the entire input sequence of length $L$.
    Let $Q=\{q_{1},q_{2},\dots,q_{|Q|}\}$ represent the token indices of the query portion in the input sequence. Then, for each document chunk $C_{i}$ (corresponding to token indices $\{c_{i,1},\dots,c_{i,K}\}$), we compute its average attention weight toward the query tokens as:
    $\text{AttentionWeight}(C_{i})=\frac{1}{K}\sum_{t\in C_{i}}\frac{1}{|Q|}\sum_{j\in Q}\mathbf{A}_{t,j}$
    This reflects how much attention the document chunk as a whole assigns to the query tokens.
  • Relevance score per chunk: For each chunk $C_{i}$, we estimate its relevance with respect to the query using an external cross-encoder model [^3] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2 which serve as an approximate ”ground-truth” relevance score:
    $\text{RelevanceScore}(C_{i})=\text{CrossEncoder}(q,\text{text}(C_{i}))$
    where $\text{text}(C_{i})$ denotes the surface text of chunk $C_{i}$.
  • Attention-Relevance Alignment Score (ARAS): For each query-document pair $(q,d)$, the ARAS is defined as the Spearman rank correlation coefficient between the attention weights and relevance scores across all chunks:
    $\text{ARAS}(q,d)=\text{Spearman}\left(\{\text{AttentionWeight}(C_{i})\}_{i=1}^{M},\ \{\text{RelevanceScore}(C_{i})\}_{i=1}^{M}\right)$
    This measures how well the ranking induced by attention weights aligns with the ranking of relevance scores across chunks. For reporting, we compute the overall mean ARAS across the samples:
    $\overline{\text{ARAS}}=\frac{1}{N}\sum_{j=1}^{N}\text{ARAS}(q_{j},d_{j})$
    This average ARAS indicates the typical alignment strength per query-document pair.
  • Positive Correlation Rate (PCR): Given a collection $\mathcal{D}=\{(q_{j},d_{j})\}_{j=1}^{N}$ of $N$ query-document pairs, we compute PCR as the proportion of pairs whose ARAS is positive:
    $\text{PCR}=\frac{1}{N}\sum_{j=1}^{N}\mathbb{I}\left(\text{ARAS}(q_{j},d_{j})>0\right)$
    where $\mathbb{I}(\cdot)$ is the indicator function. PCR reflects how consistently attention correlates positively with relevance across the dataset.
ARAS focuses on the alignment strength between attention and relevance for individual query-document pairs, while PCR reflects alignment stability across the dataset. Together, these metrics offer both instance-level and corpus-level interpretability.
3.3.2. Experimental Setup
We conduct controlled experiments on 500 query-relevant document pairs sampled from the same development set (MS MARCO) as Section 3.2. Each document is truncated to 1200 tokens, and segmented into 64-token chunks for analysis. The ARAS and PCR metrics are then computed across three representative attention heads previously identified: Layer 1 Head 23, Layer 8 Head 25, and Layer 31 Head 24.
To simulate long irrelevant contexts, we insert 800 to 1800 noise tokens either before or after the relevant document content, allowing us to test attention stability under varying noise levels.
3.3.3. Results and Findings
Layer 1, Head 23.
Under clean input, ARAS achieves 0.349 and PCR reaches 83.9%, suggesting moderate attention-relevance alignment at shallow layers. However, performance degrades significantly as noise is inserted, particularly when noise follows the relevant content. ARAS drops to negative values and PCR falls below 50% under heavy noise, indicating this head is highly sensitive to positional disruptions.
Layer 8, Head 25.
This head shows the strongest relevance alignment: baseline ARAS reaches 0.602 and PCR 96.6%. Although both metrics decline with added noise, PCR remains relatively stable (above 70%), suggesting that this mid-layer head is better at resisting irrelevant information. Nevertheless, ARAS still declines sharply as more noise accumulates, indicating that although the head consistently identifies relevant regions, the precision of its attention distribution becomes less aligned with true relevance as noise increases.
Layer 31, Head 24.
This deep-layer head exhibits relatively weak alignment, with a baseline ARAS of 0.284 and PCR of 79.6%. As with earlier heads, both metrics decline in the presence of noise. Although its performance degrades more slowly than Layer 1, it still suffers substantial drops under high noise levels—ARAS falls below 0.1 and PCR drops to approximately 56%.
Overall Observations.
Across all heads, inserting irrelevant tokens after the relevant content consistently causes more severe alignment degradation than inserting noise before, likely because autoregressive LLMs encounter relevant content later, reducing the usable attention capacity.
3.4. Summary: Implications for Long Document Retrieval
These findings provide a direct answer to RQ1, showing that while certain attention heads do exhibit relevance-focused behaviors, their ability to preserve this alignment diminishes substantially when irrelevant content accumulates. This degradation is especially pronounced when noise appears later in the sequence, likely due to the left-to-right processing nature of decoder-only LLMs such as RankLLaMA.
These results reinforce the need for explicit block selection before LLM reranking—not only to reduce computational overhead, but also to preserve attention focus, mitigate distraction from irrelevant text, and prevent attention dispersion. This insight validates the core motivation behind KeyB2: block selection remains necessary even in the era of large language models. By selecting and ranking the most relevant content blocks, we enable LLMs to concentrate attention on relevant information, thereby improving both retrieval effectiveness and computational efficiency.
(a) Insert noise document texts before relevant document
(b) Insert noise document texts after relevant document
(c) Insert noise document texts before relevant document
(d) Insert noise document texts after relevant document
(e) Insert noise document texts before relevant document
(f) Insert noise document texts after relevant document
Figure 6. Mean ARAS scores and positive correlation rates for different layers across various noise lengths and different noise insertion positions. Baseline is the score without inserting noise tokens.
4. KeyB2 Framework: Block Selection Enhanced LLM Reranking
4.1. Recap of KeyB Framework and Differences in KeyB2
The KeyB framework (Li et al., 2023a) was originally proposed to address long-document ranking challenges by applying a two-stage architecture: (1) selecting salient blocks from documents, and (2) performing relevance estimation on the selected blocks. In KeyB, block selection is conducted using either BM25 scoring or a the cross-encoder itself. The downstream reranker in KeyB is a BERT-based cross-encoder, which jointly models the query and selected document blocks within 512 tokens.
The fundamental design insight of KeyB is that long documents typically contain many irrelevant sections, and identifying relevant blocks prior to full interaction modeling can improve both ranking effectiveness and computational efficiency. In this work, we extend this block selection paradigm to modern large language models (LLMs), resulting in the KeyB2 framework. While KeyB2 retains the similar select-then-rerank architecture as KeyB, it introduces several key differences and enhancements:
  • Architectural adaptation to autoregressive LLMs: KeyB2 replaces KeyB’s bidirectional BERT reranker with decoder-only LLMs (e.g., Llama 2/3), which necessitates a different scoring mechanism. Instead of [CLS] token pooling, KeyB2 follows the RankLLaMA approach: the entire input sequence (query + selected blocks) is encoded autoregressively, and the final hidden state is passed through a linear layer to obtain the relevance score. This design enables fine-grained modeling of long input sequences while adapting to the left-to-right generation nature of LLMs.
  • Expanded block selectors: KeyB2 extends the block selection module to support three types of selectors: BM25, neural cross-encoders, and bi-encoders. This design enables flexible trade-offs between selection accuracy and computational efficiency, accommodating diverse deployment constraints and retrieval scenarios.
  • Attention-based design motivation: Our attention analysis on LLM rerankers (Section 3) reveals that feeding inputs with appended irrelevant content—can lead to significant attention dispersion and misalignment. These findings highlight the need for explicit block selection to mitigate such attention degradation, reinforcing the design rationale behind KeyB2.